OpenStudy (anonymous):

What is the extraneous solution found in solving the equation logsmall2 4x + logsmall2 (x + 1) = 3 A) -2 B) -1 C) 1 D) 2 E) -3

OpenStudy (precal):

sub it into your equation, it will be the one that is false

OpenStudy (anonymous):

how would it be the one that is false? Don't you want to find the solution?

OpenStudy (precal):

when you solve it, you will find that x will give you that as a solution, but when you test it it does not work extraneous EXTRA solution that does not work ever hear of a false positive on an exam You took an exam -say pregnancy and it came out positive but you go to the doctor and they test your blood and find out you were never pregnant the exam gave you a false positive same thing in math except we call it extraneous

OpenStudy (anonymous):

I'm not getting it it's all coming out the wrong answer

OpenStudy (phi):

the first step to answering this question is to solve for x in \[ log_2(4x) + log_2(x+1) = 3\] to solve, I would begin by using log(a)+log(b) = log(a*b) to rewrite the equation then make both sides powers to the base 2 to get rid of the log

OpenStudy (jiteshmeghwal9):

\[\log_{2}{4x}+\log_{2}{x+1} \] =>\[\log_{2}{4x(x+1)} \] =>\[\log_{2}{4x^2+4} \]

jhonyy9 (jhonyy9):

where is x next 4 ?

OpenStudy (jiteshmeghwal9):

Oops!!! sorry it's 4x.

jhonyy9 (jhonyy9):

rewrite it please

OpenStudy (jiteshmeghwal9):

\[\log_{2}{4x}+\log_{2}{x+1} \] =>\[\log_{2}{4x(x+1)} \] =>\[\log_{2}{4x^2+4x} \]

OpenStudy (phi):

zac, are you following this?

OpenStudy (anonymous):

yes and no

OpenStudy (jiteshmeghwal9):

Then, it will be \[\log_{2}{4x^2+4x}=3 \] =>\[2^3=4x^2+4x\] Now, solve it & gt ur answer:)

OpenStudy (precal):

hint you will need to use the quadratic formula to solve for x

OpenStudy (jiteshmeghwal9):

Yeah, precal is correct.

OpenStudy (phi):

so to summarize, we started with your equation \[ log_2(4x) + log_2(x+1) =3 \] we use a (very important) property of logs: we can rewrite the sum of 2 logs (of the same base) as the log of products: that means we get \[ log_2(4x(x+1))=3 \] now make both sides the exponents to base 2: \[ 2^{log_2(4x(x+1))}=2^3 \] hopefully you know the stuff on the left "undoes" the log \[ 4x(x-1)= 8 \] can you solve for x at this point?

OpenStudy (jiteshmeghwal9):

use \[-b \pm \sqrt{b^2-4ac}\over2a\]